A consumer has $100 to spend on two goods X and Y with prices $3 and $5 respectively. Derive the equation of the budget line and sketch the graph


1
Expert's answer
2021-12-24T06:54:07-0500

The equation of the budget line is:

3X + 5Y = 100.

The budget line can be depicted as a downward-sloping line, which crosses Y axis at 100/5 = 20 units and X axis at 100/3 = 33.3 units.